OpenStudy (anonymous):

Use the quadratic formula to find the roots of the equation: x^2+x+10=0

OpenStudy (anonymous):

quadratic formula \[(-b \pm \sqrt{b ^{2}-4ac})/2a\] \[ax ^{2}+bx+c\] So a=1; b=1 and c=10 So you have \[(-1\pm \sqrt{1^{2}-(4)(1)(10)})/2(1)\] This gets you\[-1/2\pm \sqrt{-39}\] Since i equals a\[\sqrt{-1}\] You have\[-1/2\pm i \sqrt{39}\]